Dont know what this is and how to do it at all. Teacher does not explain at all.
y=2x-9
x+3y=8
Solve each system by substitution

Found 2 solutions by richard1234, stanbon:
Answer by richard1234(7193)

Substitute the expression for y from the first equation into the second equation to obtain a single-variable equation. As in,

x + 3(2x - 9) = 8

Then, you can easily solve for x, then plug in x into either equation to find y.

Answer by stanbon(75887)

y=2x-9
x+3y=8
Solve each system by substitution
-----
Substitute for "y" in the 2nd equation.
Then solve for "x":
----
x + 3(2x-9) = 8
---
x + 6x-27 = 8
7x = 8+27
----
7x = 35
---
x = 5
--------
Now solve for "y"
y = 2x-9
y = 2(5)-9
---
y = 1
=====================
Solution: (5,1)
=====================
Cheers,
Stan H.
===========